OpenStudy (anonymous):

A 120-foot microwave tower sits atop a hill. From a point on level ground, an observer notes the angle of elevation to the top of the tower is 20 degrees while the angle of elevation to the bottom of the tower is 15 degrees. what is the height of the hill above level ground to the nearest foot?

OpenStudy (anonymous):

sin A/120 =sin B/b where b=120sin20/sin15 =158.5 is not the right answer.

OpenStudy (anonymous):

maybe h=52.0 ft where c=90.8083, sin35 =h/c